OpenStudy (anonymous):

4x = 3y - 7; x= 4,5,6 SHOW WORK PLEASE

OpenStudy (anonymous):

x=4 4(4)=3y-7 16=3y-7 23=3y 23/3=y x=5 4(5)=3y-7 20=3y-7 27=3y 9=y x=6 4(6)=3y-7 24=3y-7 31=3y 31/3=y